OpenStudy (anonymous):

Express sin(4theta) - sin(2theta) as a product of sine and/or cosine

OpenStudy (freckles):

\[\sin(u)-\sin(v)=2 \cos(\frac{u+v}{2})\sin(\frac{u-v}{2})\]

OpenStudy (freckles):

do you know that identity ?

OpenStudy (anonymous):

\[\sin \alpha - \sin \beta= \frac{ 1 }{ 2 } \sin \frac{ \alpha-\beta }{2 } \cos \frac{ \alpha+\beta }{ 2 }\]